# Visual Explanations of Image-Text Representations via Multi-Modal Information Bottleneck Attribution

### v1

@@ 1 Introduction
Vision-Language Pretrained Models (VLPMs) have become the de facto standard for solving a broad range of vision-language problems (Long et al., 2022). They are pre-trained on large-scale multimodal data to learn complex associations between images and text and then fine-tuned on a given downstream task. For example, the widely used CLIP model Radford et al. (2021), which uses image and text encoders trained on 400 million image-text pairs, has demonstrated remarkable performance when used for challenging vision-language tasks, such as Visual Question Answering and Visual Entailment (Shen et al., 2022).
VLPMs are highly overparameterized black-box models, enabling them to represent complex relationships in data. For example, Vision Transformers (ViTs; Dosovitskiy et al., 2021) are a state-of-the-art transformer-based vision model and are used as the image encoder of CLIP (Radford et al., 2021), containing 12 layers and 86 million parameters for ViT-Base and 24 layers and 307 million parameters for ViT-Large. Unfortunately, VLPMs like CLIP are difficult to interpret. However, model interpretability is essential in safety-critical real-world applications where VLPMs could be applied successfully, such as clinical decision-making or image captioning for the visually impaired. Improved interpretability and understanding of VLPMs would help us identify errors and unintended biases in VLP and improve the safety, reliability, and trustworthiness of VLPMs, thereby allowing us to deploy them in such safety-critical settings.
Figure 1: Example Attribution Maps For Image and Text Inputs. The red rectangles in the second and third rows show the ground-truth bounding boxes associated with the text, provided in the MS-CXR dataset (Boecking et al., 2022). Multi-modal information bottleneck (M2IB) attribution maps successfully identify relevant objects in the given image-text pairs, while other methods provide less precise localization and neglect critical features in the inputs.
@@
While existing attribution methods focus primarily on unimodal models, we propose an attribution method for VLPMs that allows us to identify critical features in image and text inputs using the information bottleneck principle (Tishby and Zaslavsky, 2015). Unlike standard unimodal attribution methods, the proposed multi-modal information bottleneck (M2IB) attribution method does not require access to ground-truth labels.
To do this, we formulate a simple multi-modal information bottleneck principle, use a variational approximation to derive a tractable optimization objective from this principle, and optimize this objective with respect to a set of attribution parameters. Using these parameters, we are able to “turn off” all irrelevant features and only keep important ones and generate attribution maps. In contrast to unimodal information bottleneck attribution (Schulz et al., 2020), we aim to find attribution parameters that maximize the likelihood of observing features of one modality given features associated with the respective other modality. We perform a qualitative and quantitative empirical evaluation and find that M2IB is able to successfully identify important features relevant to both image and text inputs. We provide an illustrative set of attribution map examples in Figure 1.
@@ 2.1 Attribution methods
Gradient-Based Attribution.   The use of gradients as a foundational component in attribution methods has been widely explored in the interpretability literature. Simonyan et al. (2014) directly use the gradient of the target with respect to inputs as the attribution score. Building on this, Grad-CAM (Gradient-weighted Class Activation Mapping; Selvaraju et al., 2016) weights convolutional layer activations by average pixel-wise gradients to identify important regions in image inputs. Integrated Gradients (Sundararajan et al., 2017) introduces two essential axioms for attribution methods— sensitivity and implementation invariance, which motivates the application of integrated gradients along paths from a baseline input to the instance under analysis.
@@
Perterbation-Based Attribution.   Perturbation approaches involve modifying input features to observe the resulting changes in model predictions. Unlike gradient-based techniques, they eliminate the need for backpropagation, allowing the model to be treated as a complete black box. However, such methods can be computationally intensive, especially for intricate model architectures, given the need to re-evaluate predictions under many perturbation conditions. The LIME (Local Interpretable Model-agnostic Explanations; Ribeiro et al., 2016) algorithm leverages perturbation in a local surrogate model. Lundberg and Lee (2017) propose SHAP (SHapley Additive exPlanations), which employs a game-theoretic approach to attribution by utilizing perturbations to compute Shapley values for each feature. Furthermore, they propose a kernel-based estimation approach for Shapley values, called KernelSHAP, drawing inspiration from local surrogate models.
@@
Attention-based Attribution.   The rise of transformers across multiple domains in machine learning has necessitated the development of specialized attribution methods tailored to these models. Using attention as an attribution method offers a straightforward approach to deciphering and illustrating the significance of various inputs in a transformer’s decision-making process. Nevertheless, solely relying on attention is inadequate as it overlooks crucial information from the value matrices and other network layers (Chefer et al., 2020). To address this issue, Chefer et al. (2021) propose a method that learns relevancy maps through a forward pass across the attention layers with contributions from each layer cumulatively forming the aggregated relevance matrices.
@@
Information-Theoretic Attribution.   Schulz et al. (2020) use information bottleneck attribution (IBA), where an information bottleneck is inserted into a layer of a trained neural network to distill the essential features for prediction. IBA is a model-agnostic method and shows impressive results on vision models including VGG-16 (Liu and Deng, 2015) and ResNet-50 He et al. (2016). Subsequently, IBA has been applied to language transformers (Jiang et al., 2020) and was shown to also outperform other methods on this task. However, IBA has thus far been focused on only one modality and has only been adopted in supervised learning. To the best of our knowledge, there is no previous work on applying the information bottleneck principle to multi-modal models like VLPMs.
@@ 2.2 Evaluation of Attribution Methods
To tackle this issue, previous studies have resorted to degradation-based metrics Chattopadhay et al. (2018); Wang et al. (2020). The underlying principle is that eliminating features with high attribution scores should diminish performance, whereas the removal of low-attributed features, often seen as noise, should potentially enhance performance. Additionally, Hooker et al. (2019) introduce ROAR (Remove and Retrain), a methodology that deliberately degrades training and validation datasets in alignment with the attribution map, thereby accounting for potential distribution shifts. Should the attribution be precise, retraining the model using these altered datasets would result in a pronounced performance decline. Rong et al. (2022) argue that mere image masking can cause a leak of information via the mask’s shape, and thus propose a Noisy Linear Imputation strategy that replaces pixels with the average of their neighbors.
It is important to note that a visually appealing saliency map does not guarantee the efficacy of the underlying attribution method. For example, an edge detector might generate a seemingly plausible saliency map, yet it does not qualify as an attribution method because it is independent of the model under analysis. Adebayo et al. (2018) propose a sanity check designed to assess whether the outputs of attribution methods genuinely reflect the properties of the specific model under evaluation. Notably, several prominent models, including Guided Backprop (Springenberg et al., 2014) and its variants appear insensitive to model weights and consequently do not pass the sanity check.
@@ 3.1 The Information Bottleneck Principle
The information bottleneck principle (Tishby et al., 2000) provides a framework for finding compressed representations of neural network models. To obtain latent representations that reflect the most relevant information of the input data, the information bottleneck principle seeks to find a stochastic latent representation $Z$ of the input source $X$ defined by a parametric encoder $p_{Z\,|\,X}(z\,|\,x;\theta)$ that is maximally informative about a target $Y$ while constraining the mutual information between the latent representation $Z$ and the input $X$. For a representation parameterized by parameters $\theta$, this principle can be expressed as the optimization problem
@@
where $\beta$ is a Lagrange multiplier that trades off learning a latent representation that is maximally informative about the target $Y$ with learning a representation that is maximally compressive about the input $X$ (Alemi et al., 2017).
@@ 3.2 A Multi-Modal Information Bottleneck Principle
Unfortunately, for VLPMs, the loss function above is not suitable since we wish to learn interpretable latent representations using only text and vision inputs without relying on task-specific targets $Y$ that may not be available or are expensive to obtain. To formulate a multi-modal information bottleneck principle for VLPMs, we need to develop an optimization objective that is more akin to optimization objectives for self-supervised methods for image-text representation learning that only use (text, image) pairs (Radford et al., 2021; Mu et al., 2021).
This learning problem fundamentally differs from supervised attribution map learning for unimodal tasks. For example, we may have an image of a bear, $X_{\text{bear}}$, and a corresponding label, $Y_{\text{bear}}=\mathrm{``bear"}$. For a unimodal classification task, we can simply maximize $I(Y_{\text{bear}};Z_{\mathrm{bear}};\theta)-\beta I(X_{\text{bear}};Z_{\mathrm{bear}};\theta)$ with respect to $\theta$, where $Z_{\mathrm{bear}}$ is the latent representation of $X_{\mathrm{bear}}$.
In contrast, in image-text representation learning, we typically have text descriptions, such as “This is a picture of a bear” ($L^{\prime}_{\text{bear}}$) instead of labels (Radford et al., 2021). In this setting, both $X_{\text{bear}}$ and $L^{\prime}_{\text{bear}}$ are “inputs” without a pre-defined corresponding label. To obtain a task-agnostic image-text representation independent from any task-specific ground-truth labels, we would like to use both input modalities and define a multi-modal information bottleneck principle and whereas the outputs are closely dependent on the specific downstream task. This requires defining an alternative to the “fitting term” $I(Z,Y;\theta)$ of the conventional information bottleneck objective.
@@
where $\𝑚m^{\prime}=\mathcal{M}\backslash m$ is the complement of $m$, and $E_{m^{\prime}}$ is embedding of modality $m^{\prime}$. Next, we will show how to use this multi-modal variant of the information bottleneck principle for attribution.
@@ 3.3 A Multi-Modal Information Bottleneck for Attribution
To restrict the information flow in the latent representation with a simple parametric encoder, we adapt the masking approach in Schulz et al. (2020). For clarity and brevity, we henceforth represent the $V_{m}\times W_{m}$-dimensional latent representation $Z_{m}$ in its vectorized form in $\mathbb{R}^{J}$ with $J\,\dot{=}\,V_{m}\cdot W_{m}$. Assuming independence across latent representation dimensions, we then define
@@
$\mathcal{F}_{\mathrm{image}}(\theta_{\mathrm{image}})$ $=I(Z_{\mathrm{image}},E_{\mathrm{text}};\theta_{\mathrm{image}})-\beta_{\mathrm{image}}\,I(Z_{\mathrm{image}},X_{\mathrm{text}};\theta_{\mathrm{image}})$ (6)
$\mathcal{F}_{\mathrm{text}}(\theta_{\mathrm{text}})$ $=I(Z_{\mathrm{text}},E_{\mathrm{image}};\theta_{\mathrm{text}})-\beta_{\mathrm{text}}\,I(Z_{\mathrm{text}},X_{\mathrm{image}};\theta_{\mathrm{text}}),$ (7)
@@ 3.4 A Variational Objective for Multi-Modal Information Bottleneck Attribution
The derivation of this objective has closely followed the steps in (Alemi et al., 2017). In practice, the objective can be computed using the empirical data distribution so that
@@ 4 Empirical Evaluation
We evaluate the proposed attribution method using CLIP (Radford et al., 2021) on four image-caption datasets, including widely-used image captioning datasets and medical datasets. Our main datasets are (i) Conceptual Captions (Sharma et al., 2018) consisting of diverse images and captions from the web, and (ii) MS-CXR (Local Alignment Chest X-ray dataset; Boecking et al. (2022)), which contains chest X-rays and texts describing radiological findings, complementing MIMIC-CXR (MIMIC Chest X-ray; Johnson et al. (2019)) by improving the bounding boxes and captions. In addition, we also include some qualitative examples from the following radiology and remote sensory datasets to show the potential application of our model in safety-critical domains. Namely, we have (iii) ROCO (Radiology Objects in COntext; Pelka et al. (2018)) that includes radiology image-caption pairs from the open-access biomedical literature database PubMed Central, and (iv) RSICD (Remote Sensing Image Captioning Dataset; Lu et al. (2017)), which collects remote sensing images from web map services including Google Earth and provides corresponding captions.
@@ 4.1 Experiment Setup
For all experiments, we use a pretrained CLIP model with ViT-B/32 (Dosovitskiy et al., 2021) as the image encoder and a 12-layer self-attention transformer as the text encoder. For Conceptual Captions, we use the pretrained weights of openai/clip-vit-base-patch32.111Cf. https://huggingface.co/openai/clip-vit-base-patch32. For MS-CXR, we use CXR-RePaiR (Endo et al., 2021) which is CLIP finetuned on radiology datasets, and compare the impact of finetuning in Section 4.5. For each {image, caption} pair, we insert an information bottleneck into the given layer of the text encoder and image encoder of CLIP separately, then train the bottleneck using the same setup as the Per-Sample Bottleneck of original IBA (Schulz et al., 2020), which duplicates a single sample for 10 times to stabilize training and runs 10 iterations using the Adam optimizer with a learning rate of 1. Experiments show no significant difference between different learning rates and more training steps. We conduct a hyper-parameter tuning on the index of the layer $l$, the scaling factor $\beta$, and the variance $\sigma^{2}$. For a discussion of these hyperparameters in the multi-modal information bottleneck objective, see Appendix A.
@@ 4.2 Qualitative Results
We qualitatively compare our method with 5 widely used attribution methods, including gradient-based GradCAM (Selvaraju et al., 2016) and Saliency method (Simonyan et al., 2014), perturbation-based Kernel SHAP Lundberg and Lee (2017) and RISE Petsiuk et al. (2018), and transformer-specific method Chefer et al. (2021). We show one image-caption example and two radiology examples in Figure 1 and provide more qualitative comparisons in the Appendix B. As shown, our method is able to capture all relevant objects appearing in both modalities, while other methods tend to focus on one major object. As illustrated in Figure 1(a), the highlighted areas in the image change according to different inputs, and our method can capture complicated relationships between image and text when involving multiple objects. Our method is also able to detect multiple occurrences of the same object in the image, as illustrated in Figure 1(b).

### v2

@@ 1 Introduction
Vision-Language Pretrained Models (VLPMs) have become the de facto standard for solving a broad range of vision-language problems [15]. They are pre-trained on large-scale multimodal data to learn complex associations between images and text and then fine-tuned on a given downstream task. For example, the widely used CLIP model [21], which uses image and text encoders trained on 400 million image-text pairs, has demonstrated remarkable performance when used for challenging vision-language tasks, such as Visual Question Answering and Visual Entailment [27].
VLPMs are highly overparameterized black-box models, enabling them to represent complex relationships in data. For example, Vision Transformers [ViTs; 7] are a state-of-the-art transformer-based vision model and are used as the image encoder of CLIP [21], containing 12 layers and 86 million parameters for ViT-Base and 24 layers and 307 million parameters for ViT-Large. Unfortunately, VLPMs like CLIP are difficult to interpret. However, model interpretability is essential in safety-critical real-world applications where VLPMs could be applied successfully, such as clinical decision-making or image captioning for the visually impaired. Improved interpretability and understanding of VLPMs would help us identify errors and unintended biases in VLP and improve the safety, reliability, and trustworthiness of VLPMs, thereby allowing us to deploy them in such safety-critical settings.
Figure 1: Example Attribution Maps For Image and Text Inputs. The red rectangles in the second and third rows show the ground-truth bounding boxes associated with the text, provided in the MS-CXR dataset [3]. Multi-modal information bottleneck (M2IB) attribution maps successfully identify relevant objects in the given image-text pairs, while other methods provide less precise localization and neglect critical features in the inputs.
@@
While existing attribution methods focus primarily on unimodal models, we propose an attribution method for VLPMs that allows us to identify critical features in image and text inputs using the information bottleneck principle [31]. Unlike standard unimodal attribution methods, the proposed multi-modal information bottleneck (M2IB) attribution method does not require access to ground-truth labels.
To do this, we formulate a simple multi-modal information bottleneck principle, use a variational approximation to derive a tractable optimization objective from this principle, and optimize this objective with respect to a set of attribution parameters. Using these parameters, we are able to “turn off” all irrelevant features and only keep important ones and generate attribution maps. In contrast to unimodal information bottleneck attribution [24], we aim to find attribution parameters that maximize the likelihood of observing features of one modality given features associated with the respective other modality. We perform a qualitative and quantitative empirical evaluation and find that M2IB is able to successfully identify important features relevant to both image and text inputs. We provide an illustrative set of attribution map examples in Figure 1.
@@ 2.1 Attribution methods
Gradient-Based Attribution.   The use of gradients as a foundational component in attribution methods has been widely explored in the interpretability literature. Simonyan et al. [28] directly use the gradient of the target with respect to inputs as the attribution score. Building on this, Grad-CAM [Gradient-weighted Class Activation Mapping; 25] weights convolutional layer activations by average pixel-wise gradients to identify important regions in image inputs. Integrated Gradients [30] introduces two essential axioms for attribution methods— sensitivity and implementation invariance, which motivates the application of integrated gradients along paths from a baseline input to the instance under analysis.
@@
Perterbation-Based Attribution.   Perturbation approaches involve modifying input features to observe the resulting changes in model predictions. Unlike gradient-based techniques, they eliminate the need for backpropagation, allowing the model to be treated as a complete black box. However, such methods can be computationally intensive, especially for intricate model architectures, given the need to re-evaluate predictions under many perturbation conditions. The LIME [Local Interpretable Model-agnostic Explanations; 22] algorithm leverages perturbation in a local surrogate model. Lundberg and Lee [17] propose SHAP (SHapley Additive exPlanations), which employs a game-theoretic approach to attribution by utilizing perturbations to compute Shapley values for each feature. Furthermore, they propose a kernel-based estimation approach for Shapley values, called KernelSHAP, drawing inspiration from local surrogate models.
@@
Attention-based Attribution.   The rise of transformers across multiple domains in machine learning has necessitated the development of specialized attribution methods tailored to these models. Using attention as an attribution method offers a straightforward approach to deciphering and illustrating the significance of various inputs in a transformer’s decision-making process. Nevertheless, solely relying on attention is inadequate as it overlooks crucial information from the value matrices and other network layers [5]. To address this issue, Chefer et al. [6] propose a method that learns relevancy maps through a forward pass across the attention layers with contributions from each layer cumulatively forming the aggregated relevance matrices.
@@
Information-Theoretic Attribution.   Schulz et al. [24] use information bottleneck attribution (IBA), where an information bottleneck is inserted into a layer of a trained neural network to distill the essential features for prediction. IBA is a model-agnostic method and shows impressive results on vision models including VGG-16 [14] and ResNet-50 [10]. Subsequently, IBA has been applied to language transformers [12] and was shown to also outperform other methods on this task. However, IBA has thus far been focused on only one modality and has only been adopted in supervised learning. To the best of our knowledge, there is no previous work on applying the information bottleneck principle to multi-modal models like VLPMs.
@@ 2.2 Evaluation of Attribution Methods
To tackle this issue, previous studies have resorted to degradation-based metrics [4, 33]. The underlying principle is that eliminating features with high attribution scores should diminish performance, whereas the removal of low-attributed features, often seen as noise, should potentially enhance performance. Additionally, Hooker et al. [11] introduce ROAR (Remove and Retrain), a methodology that deliberately degrades training and validation datasets in alignment with the attribution map, thereby accounting for potential distribution shifts. Should the attribution be precise, retraining the model using these altered datasets would result in a pronounced performance decline. Rong et al. [23] argue that mere image masking can cause a leak of information via the mask’s shape, and thus propose a Noisy Linear Imputation strategy that replaces pixels with the average of their neighbors.
It is important to note that a visually appealing saliency map does not guarantee the efficacy of the underlying attribution method. For example, an edge detector might generate a seemingly plausible saliency map, yet it does not qualify as an attribution method because it is independent of the model under analysis. Adebayo et al. [1] propose a sanity check designed to assess whether the outputs of attribution methods genuinely reflect the properties of the specific model under evaluation. Notably, several prominent models, including Guided Backprop [29] and its variants appear insensitive to model weights and consequently do not pass the sanity check.
@@ 3.1 The Information Bottleneck Principle
The information bottleneck principle [32] provides a framework for finding compressed representations of neural network models. To obtain latent representations that reflect the most relevant information of the input data, the information bottleneck principle seeks to find a stochastic latent representation $Z$ of the input source $X$ defined by a parametric encoder $p_{Z\,|\,X}(z\,|\,x;\theta)$ that is maximally informative about a target $Y$ while constraining the mutual information between the latent representation $Z$ and the input $X$. For a representation parameterized by parameters $\theta$, this principle can be expressed as the optimization problem
@@
where $\beta$ is a Lagrange multiplier that trades off learning a latent representation that is maximally informative about the target $Y$ with learning a representation that is maximally compressive about the input $X$ [2].
@@ 3.2 A Multi-Modal Information Bottleneck Principle
Unfortunately, for VLPMs, the loss function above is not suitable since we wish to learn interpretable latent representations using only text and vision inputs without relying on task-specific targets $Y$ that may not be available or are expensive to obtain. To formulate a multi-modal information bottleneck principle for VLPMs, we need to develop an optimization objective that is more akin to optimization objectives for self-supervised methods for image-text representation learning that only use (text, image) pairs [18, 21].
This learning problem fundamentally differs from supervised attribution map learning for unimodal tasks. For example, we may have an image of a bear, $X_{\text{bear}}$, and a corresponding label, $Y_{\text{bear}}=\mathrm{``bear"}$. For a unimodal classification task, we can simply maximize $I(Y_{\text{bear}},Z_{\mathrm{bear}};\theta)-\beta I(X_{\text{bear}},Z_{\mathrm{bear}};\theta)$ with respect to $\theta$, where $Z_{\mathrm{bear}}$ is the latent representation of $X_{\mathrm{bear}}$.
In contrast, in image-text representation learning, we typically have text descriptions, such as “This is a picture of a bear” ($L^{\prime}_{\text{bear}}$) instead of labels [21]. In this setting, both $X_{\text{bear}}$ and $L^{\prime}_{\text{bear}}$ are “inputs” without a pre-defined corresponding label. To obtain a task-agnostic image-text representation independent from any task-specific ground-truth labels, we would like to use both input modalities and define a multi-modal information bottleneck principle and whereas the outputs are closely dependent on the specific downstream task. This requires defining an alternative to the “fitting term” $I(Z,Y;\theta)$ of the conventional information bottleneck objective.
@@
where $\𝑚m^{\prime}=\mathcal{M}\backslash m$ is the complement of $m$, and $E_{m^{\prime}}$ is the embedding of modality $m^{\prime}$. Next, we will show how to use this multi-modal variant of the information bottleneck principle for attribution.
@@ 3.3 A Multi-Modal Information Bottleneck for Attribution
To restrict the information flow in the latent representation with a simple parametric encoder, we adapt the masking approach in Schulz et al. [24]. For clarity and brevity, we henceforth represent the $V_{m}\times W_{m}$-dimensional latent representation $Z_{m}$ in its vectorized form in $\mathbb{R}^{J}$ with $J\,\dot{=}\,V_{m}\cdot W_{m}$. Assuming independence across latent representation dimensions, we then define
@@
$\mathcal{F}_{\mathrm{image}}(\theta_{\mathrm{image}})$ $=I(Z_{\mathrm{image}},E_{\mathrm{text}};\theta_{\mathrm{image}})-\beta_{\mathrm{image}}\,I(Z_{\mathrm{image}},X_{\mathrm{image}};\theta_{\mathrm{image}})$ (6)
$\mathcal{F}_{\mathrm{text}}(\theta_{\mathrm{text}})$ $=I(Z_{\mathrm{text}},E_{\mathrm{image}};\theta_{\mathrm{text}})-\beta_{\mathrm{text}}\,I(Z_{\mathrm{text}},X_{\mathrm{text}};\theta_{\mathrm{text}}),$ (7)
@@ 3.4 A Variational Objective for Multi-Modal Information Bottleneck Attribution
The derivation of this objective has closely followed the steps in [2]. In practice, the objective can be computed using the empirical data distribution so that
@@ 4 Empirical Evaluation
We evaluate the proposed attribution method using CLIP [21] on four image-caption datasets, including widely-used image captioning datasets and medical datasets. Our main datasets are (i) Conceptual Captions [26] consisting of diverse images and captions from the web, and (ii) MS-CXR (Local Alignment Chest X-ray dataset; [3]), which contains chest X-rays and texts describing radiological findings, complementing MIMIC-CXR (MIMIC Chest X-ray; [13]) by improving the bounding boxes and captions. In addition, we also include some qualitative examples from the following radiology and remote sensory datasets to show the potential application of our model in safety-critical domains. Namely, we have (iii) ROCO (Radiology Objects in COntext; Pelka et al. [19]) that includes radiology image-caption pairs from the open-access biomedical literature database PubMed Central, and (iv) RSICD (Remote Sensing Image Captioning Dataset; Lu et al. [16]), which collects remote sensing images from web map services including Google Earth and provides corresponding captions.
@@ 4.1 Experiment Setup
For all experiments, we use a pretrained CLIP model with ViT-B/32 [7] as the image encoder and a 12-layer self-attention transformer as the text encoder. For Conceptual Captions, we use the pretrained weights of openai/clip-vit-base-patch32.111Cf. https://huggingface.co/openai/clip-vit-base-patch32. For MS-CXR, we use CXR-RePaiR [8] which is CLIP finetuned on radiology datasets, and compare the impact of finetuning in Section 4.5. For each {image, caption} pair, we insert an information bottleneck into the given layer of the text encoder and image encoder of CLIP separately, then train the bottleneck using the same setup as the Per-Sample Bottleneck of original IBA [24], which duplicates a single sample for 10 times to stabilize training and runs 10 iterations using the Adam optimizer with a learning rate of 1. Experiments show no significant difference between different learning rates and more training steps. We conduct a hyper-parameter tuning on the index of the layer $l$, the scaling factor $\beta$, and the variance $\sigma^{2}$. For a discussion of these hyperparameters in the multi-modal information bottleneck objective, see Appendix A.
@@ 4.2 Qualitative Results
We qualitatively compare our method with 5 widely used attribution methods, including gradient-based GradCAM [25] and Saliency method [28], perturbation-based Kernel SHAP [17] and RISE [20], and transformer-specific method [6]. We show one image-caption example and two radiology examples in Figure 1 and provide more qualitative comparisons in the Appendix B. As shown, our method is able to capture all relevant objects appearing in both modalities, while other methods tend to focus on one major object. As illustrated in Figure 2(a), the highlighted areas in the image change according to different inputs, and our method can capture complicated relationships between image and text when involving multiple objects. Our method is also able to detect multiple occurrences of the same object in the image, as illustrated in Figure 2(b).